OpenStudy (anonymous):

Factor completely: -5x^2 + 10x - 15

OpenStudy (anonymous):

\[ -5x^2 + 10x - 15 = -5 (x^2 - 2x + 3). \]Notice that the quadratic cannot be factored into reals. You can see this by finding the roots via the quadratic formula. In fact, you see that the discriminant is 2^2 - 12, which is negative. Thus, we cannot factor any further than \[ \boxed{-5(x^2-2x+3)}. \]

OpenStudy (calculusfunctions):

-5x² + 10x − 15 First the GCF is -5, hence -5x² + 10x − 15 = -5(x² − 2x + 3) There are no two integers which yield a product of +3 and a sum of -2. Therefore it cannot be factored over integers.

OpenStudy (anonymous):

Thank you soo much :)

OpenStudy (anonymous):

No problem. Thanks for the medal! :)